# Integrating Protein Sequence and Expression Level to Analysis Molecular Characterization of Breast Cancer Subtypes

### v1

@@ 5 Conclusion
This research presents an innovative approach to understanding breast cancer by integrating protein sequence and expression data using ProtGPT2 and machine learning techniques such as XGBoost and Random Forest to enhance subtype classification and predict clinical outcomes. By fusing structural protein information with quantitative expression levels, this study creates a comprehensive representation of the proteins involved, improving accuracy in predicting key biomarkers (ER, PR, HER2) and enabling the clustering of patients into meaningful subgroups. The use of ensemble clustering further highlights the heterogeneity of breast cancer, revealing distinct molecular patterns that align with known subtypes and suggesting targeted treatment opportunities. Additionally, protein-protein interaction analysis reveals meaningful functional relationships, such as those involving proteins like KMT2C and CLASP2, which play critical roles in tumor progression and could serve as therapeutic targets. This integrated approach demonstrates the potential of combining cutting-edge proteomic analysis with machine learning to advance precision oncology, offering a framework that could be extended to other complex diseases for deeper biological insights and more effective personalized treatments.

### v2

@@ 5 Conclusion
This study introduces a novel framework for breast cancer analysis by integrating protein sequence with expression data using ProtGPT2-derived embeddings, and enhanced with machine learning models such as XGBoost for subtype classification and predict clinical outcomes. By combining structural protein information with quantitative expression levels, this study creates a comprehensive representation of the proteins involved, improving accuracy in predicting key biomarkers (ER, PR, HER2) and enabling the clustering of patients into meaningful subgroups. Ensemble clustering revealed the molecular heterogeneity of breast cancer, uncovering patterns aligned with known subtypes and suggesting avenues for targeted therapies. Additionally, protein–protein interaction analysis reveals meaningful functional relationships, such as those involving proteins like KMT2C, linked to estrogen receptor regulation, and MYO1B, which contributes to drug resistance, and invasion in hormone receptor-positive and triple-negative breast cancers, highlighting their potential as therapeutic targets within a biologically informed framework. Overall, this integrated approach demonstrates the power of combining advanced protein language models with expression-informed embeddings, enhanced by machine learning, to advance precision oncology and provide a framework that can be extended to other complex diseases for deeper biological insights and more effective personalized treatments.
Appendix A Supplementary Figures
(a) Co-occurrence matrix of the Orange Cluster
(b) Co-occurrence matrix of the Green Cluster
(c) Co-occurrence matrix of the Red Cluster
Figure A.1: Co-occurrence matrices for three patient clusters (Orange, Green, and Red), illustrating correlations between breast cancer markers.